OpenStudy (thickgirl_angel):

Complete the square on the following quadratic x^2+22x+9=0 what value is added to both sides of the equation.

OpenStudy (imstuck):

I don't understand this question. Did you do something to it when you posted it or is that the way it was presented to you? What exactly are you trying to do with that?

OpenStudy (imstuck):

The original problem is x^2 + 22x + 9 = 0?

OpenStudy (thickgirl_angel):

Yes

OpenStudy (imstuck):

Do you know how to complete the square?

OpenStudy (thickgirl_angel):

I use too

OpenStudy (imstuck):

Ok, there are steps you follow every time to complete the square. I'll type them here and then we can do the problem...

OpenStudy (imstuck):

Separate the x terms from the other ones with the equal sign. Make sure the coefficient on the x^2 term is 1. Take half the x term, square it, and add it to both sides. Let's do those first.

OpenStudy (imstuck):

\[x ^{2}+2x+9=0\] The first step is to make sure the coefficient on the x^2 is 1 and it is, so we can move on.

OpenStudy (imstuck):

Separate the x terms from everything else by moving everything else to the other side of the equals sign. \[x ^{2}+22x=-9\]

OpenStudy (imstuck):

What is the coefficient on the x term?

OpenStudy (thickgirl_angel):

-9?

OpenStudy (imstuck):

The coefficient on the x term is 22. See up above? 22x

OpenStudy (imstuck):

Sorry I was gone for a minute...

OpenStudy (imstuck):

We need to take half the value of that x term; take half the value of 22 and what do we have? What is half of 22?

OpenStudy (thickgirl_angel):

11

OpenStudy (imstuck):

Ok, next step is to square that number. What is 11 times 11?

OpenStudy (thickgirl_angel):

121

OpenStudy (imstuck):

Yes! THAT is the number you add to both sides of the equation. This is what you have then: \[x ^{2}+22x+121=-9+121\]

OpenStudy (imstuck):

That is the critical point where you just completed the square, because what you have done on the left side is created a perfect square trinomial that can be factored into this: \[(x+11)^{2}\]

OpenStudy (imstuck):

How far do you need to go with this? Do you have to solve it for x or what?

OpenStudy (thickgirl_angel):

Just what is added to both side

OpenStudy (imstuck):

ok then!

OpenStudy (thickgirl_angel):

I Like when you help me

OpenStudy (imstuck):

Thank you! That means a lot!

OpenStudy (thickgirl_angel):

Your Welcome :)